# CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift

### v1

@@
Massive, intensely star-forming galaxies at high redshift require a supply of molecular gas from their gas reservoirs, replenished by infall from the surrounding circumgalactic medium, to sustain their immense star-formation rates. However, our knowledge of the extent and morphology of their cold-gas reservoirs is still in its infancy. We present the results of stacking 80 hours of JVLA observations of CO(1–0) emission – which traces the cold molecular gas – in 19 dusty, star-forming galaxies from the AS2VLA survey. The visibility-plane stack reveals extended emission with a half-light radius of kpc, 2–3 more extended than the dust-obscured star formation and more extended than the stellar emission. Similarly, stacking the [C i](1–0) observations for a subsample of our galaxies yields sizes consistent with CO(1–0). The CO(1–0) size is comparable to the [C ii] halos detected around high-redshift star-forming galaxies. The bulk (up to 80%) of molecular gas resides outside the star-forming region; only a small part of their molecular gas reservoir directly contributes to their current star formation. Photon-dissociation region modelling indicates that the extended CO(1–0) emission arises from clumpy, dense clouds rather than smooth, diffuse gas.
@@
The majority of star formation beyond redshift took place in massive, dusty star-forming galaxies (DSFGs111Also know as sub-millimetre galaxies – SMGs.) with star-formation rates (SFRs) of a few 100 – 1000 M⊙ yr-1 (e.g., Magnelli et al. 2011; Dudzevičiūtė et al. 2020; Zavala et al. 2021). While some DSFGs are starbursts triggered by major mergers, many appear to be relatively isolated (Gillman et al., 2024; Hodge et al., 2024). The immense SFRs of these isolated galaxies need to be fed by a continuous supply of cold gas from their extended gas reservoirs and the surrounding circumgalactic medium (CGM, e.g., Tumlinson et al. 2017; Faucher-Giguère & Oh 2023) and by minor mergers with gas-rich satellites (e.g., Narayanan et al. 2015).
@@
In local galaxies, cold, molecular gas is primarily traced by the CO(1–0) rotational line, which has an excitation temperature of just 5.5 K and critical density of cm-3. At high redshift, studies of molecular gas have mostly focused on higher-excitation CO rotational emission lines (see, e.g., Hodge & da Cunha 2020; Tacconi et al. 2020; Walter et al. 2020). However, the mid/high-excitation CO emission lines arise from the denser, warmer gas, and are known to be much more compact than the ground-state CO(1–0) transition (e.g., Leroy et al. 2009; Ivison et al. 2011; Emonts et al. 2016), making them unsuitable for tracing cold, diffuse gas.
@@
In contrast to the higher-excitation CO lines, studies of the ground-state CO(1–0) line have been scarce and largely focused on the few brightest galaxies. This is because of the low intrinsic brightness of the CO(1–0) line, a large collecting area and long integrating times are required for achieving the necessary signal-to-noise ratio.
@@
The CO(1–0) emission in a number of high-redshift galaxies has been observed by the Green Bank Telescope (e.g., Hainline et al. 2006; Swinbank et al. 2010; Frayer et al. 2011; Harris et al. 2012; Harrington et al. 2021). However, these spatially unresolved observations do not provide any information about the source size. Early Karl. G. Jansky Very Large Array (JVLA) studies of CO(1–0) emission in bright high-redshift DSFGs (e.g., Ivison et al. 2010, 2011; Riechers et al. 2011b; Thomson et al. 2012) reported half-light radii ranging from a few kpc to 25 kpc (J02399-0136, Thomson et al. 2012), often showing multiple CO(1–0) components. However, some of these results have been challenged by deeper observations: for example, Frias Castillo et al. (2022) showed that the supposedly extended CO(1–0) reservoir with high-velocity wings in the galaxy J13120 (Riechers et al., 2011a) is likely an artefact of low signal-to-noise of the original data. To summarise, the sizes of CO(1–0) reservoirs reported by these early studies might be biased towards the brightest (i.e., most extreme) sources, and suffer from low S/N.
@@
More recently, apparently very extended molecular gas reservoirs have been reported in high-density environments of early protoclusters, such as the Spiderweb (, Emonts et al. 2016) which shows a CO(1–0) reservoirs with a half-light radius of kpc CO(1–0), the -kpc CO(4–3) emission around the star-forming galaxy from Ginolfi et al. (2017), or the claimed super-extended, 200 kpc CO(3–2) emission around a quasar cid_346 (Cicone et al. 2021, but c.f. non-detection in deeper data by Jones et al. 2023).
@@
Further evidence for extended gas around high-redshift DSFGs comes from Atacama Large (Sub)Millimetre Array (ALMA) observations of the [C ii] 158-m line. Extended [C ii] “halos” have been observed around high-redshift DSFGs (e.g., Gullberg et al. 2018; Rybak et al. 2019, 2020) as well as the less massive star-forming galaxies (e.g., Fujimoto et al. 2019, 2020; Ginolfi et al. 2020), sometimes on scales as large as 50 kpc.
@@
The ground-state CO(1–0) emission line remains the most robust probe of diffuse molecular gas. In the recent years, there have been systematic surveys of CO(1–0) in high-redshift DSFGs using JVLA (Frias Castillo et al., 2023; Stanley et al., 2023). These have yielded tentative evidence for extended CO(1–0) emission: Frias Castillo et al. (2023) found that the CO(1–0) extracted from moment-0 maps increases with the aperture size out to radii of 5 (40 kpc). Similarly, using 1-resolution JVLA imaging of 14 galaxies from the z-GAL survey (Cox et al., 2023), Stanley et al. (2023) found CO(1–0) reservoirs of up to 25 kpc across. However, it is unclear how many of their sources are magnified by gravitational lensing, making the interpretation of the measured source sizes difficult.
In this Paper, we report results of stacking 80 hours of CO(1–0) observations from the recently completed AS2VLA222AS2VLA = ALMA, SCUBA-2, VLA. survey (Frias Castillo et al., 2023). Specifically, by using JVLA in the most compact D-array, we combined high surface-brightness sensitivity with ability to recover extended emission. This setup makes them ideally suited for studying potentially very extended, low surface-brightness reservoirs around DSFGs. We identify extended CO(1–0) reservoirs with a scale radius of 4 kpc, indicating that in high-redshift galaxies, the molecular gas extends well beyond the dust and (potentially) stellar emission. The CO(1–0)/[C ii] and dust continuum surface brightness ratios in these reservoirs are consistent with in-situ star-formation, indicating that the extended CO(1–0) and [C ii] halos are a natural extension of the galaxy’s gas reservoirs, rather than residuals of galactic outflows.
@@ 2.1 Target sample
Our initial sample consists of 30 sub-mm selected, star-forming galaxies at from the AS2VLA survey (Frias Castillo et al., 2023, Paper I). The sample is selected from ALMA continuum detections of source taken from 4 deg2 of SCUBA-2 850m maps from the S2CLS (Geach et al., 2017) and S2COSMOS (Simpson et al., 2019) surveys. Individual sources were pin-pointed by high-resolution ALMA and SMA imaging (Stach et al., 2018; Hill et al., 2018; Simpson et al., 2020); the 870-m fluxes range between 7.5 and 18.3 mJy. All sources have spectroscopic redshifts based on mid- CO lines and [C i](1–0) (Birkin et al. 2021; Chen et al. 2022a, b, Chapman et al., in prep.). The sources span stellar mass M⊙ and SFR = 400 – 1700 M⊙ yr-1, with a median SFR of 1010 M⊙ yr-1 (Frias Castillo et al., 2023; Liao et al., 2024). Most of them are on the massive end of the main sequence of star-forming galaxies, but several are offset to the starburst regime (SFR/ higher than the main-sequence, Frias Castillo et al. 2023). None of the galaxies is strongly gravitationally lensed; we can thus directly measure their intrinsic sizes.
@@ 2.2 JVLA observations
We used JVLA to observe the CO(1–0) emission (rest-frame frequency: = 115.2712 GHz) in thirty DSFGs at . The surveys comprises of the JVLA programs ID:21A-254 (PI: Hodge) and ID:17A-251 (PI: Walter). The observations were taken between March 2021 and December 2023, under good weather conditions in the D-array configuration. The total time per source including overheads was four hours for almost all sources. The exceptions were AS2COS14.1, which was observed for eight hours, and AS2COS44.1 and AS2COS65.1, which were observed for two hours each.
We observed in the K- and Ka-bands, using a contiguous bandwidth of 2 GHz with 2-MHz spectral channels. Depending on the source redshift, the angular resolution ranged from 2.5 to 5.1 (circularised FWHM); the maximum recoverable scale was 44–60. We manually inspected and calibrated the data using the Common Astronomy Software Applications package (Casa, McMullin et al. 2007) versions 6.4.1 and 6.4.3 and the standard JVLA pipeline. The observations of the first 17 galaxies (eleven secure detections) were presented in Frias Castillo et al. (2023). In this work (Paper II), we focus on the 19 sources with secure CO(1–0) detections333For a “secure” detection, we require peak S/N3 in naturally-weighted cleaned maps of CO(1–0) emission. from the completed survey (Paper III, in prep).
@@ 2.3 Visibility-plane analysis and stacking
We measure the source sizes directly in the visibility-plane. This approach bypasses several issues with the Cleaned images, including varying beam shape, the fact that cleaned images combine the dirty and clean (Gaussian) beam, and the correlated noise in the image-plane. For illustration, we show the stacked moment-0 image in Fig. 1. Image-plane deconvolution with the imfit task in Casa does not yield a reliable source size.
@@
For the visibility-plane analysis, we process the data as follows: (1) if the wide-band imaging shows a significant (S/N5) continuum, subtract the continuum (constant signal) from the data using the uvcontsub task; (2) select the velocity range within 1 FWHM from the systemic velocity (using linewidths from mid- CO lines, which are detected at a higher S/N); (3) frequency-average the selected data into a single channel; (4) bin the resulting visibilities in radial bins with a size of 2.5, 5, and 10 k. The fields were centred at the ALMA dust continuum peak (accurate to 0.1). We confirmed that the phase-tracking centre is well-aligned with the CO(1–0) emission; we phase-shifted three sources with noticeable offsets (AEG2, CDFN1, CDFN2). For two sources (AS2COS14.1, AS2COS31.1), we found a bright continuum source within the VLA primary beam. We subtracted these companions from the visibilities in Casa, using the ft task.
@@
We fit each binned visibility function with: (1) point-source (constant signal) and (2) an exponential profile. We prefer the exponential profile to the more widely-used Gaussian profile as it provides a more physical description of gas surface brightness distribution in star-forming galaxies 444For the literature source sizes derived using Gaussian profiles, we convert the half-light radius from the Gaussian to exponential profile. (e.g., Leroy et al. 2009; Wang & Lilly 2022) as well as dust and stellar light distribution (e.g., Hodge et al. 2016; Gullberg et al. 2019; Gillman et al. 2024. We utilise the Monte-Carlo Markov chain modelling using the emcee package (Foreman-Mackey et al., 2013). We assumed wide, uniform priors in total flux and half-light radius. For each galaxy, we calculate the Akaike and Bayesian information criteria (AIC, BIC) for each model. The results for individual galaxies are listed in Table 2.

### v2

@@
11email: mrybak@strw.leidenuniv.nl 22institutetext: Faculty of Electrical Engineering, Mathematics and Computer Science, Delft University of Technology, Mekelweg 4, 2628 CD Delft, The Netherlands 33institutetext: SRON – Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands 44institutetext: Centre for Extragalactic Astronomy, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 55institutetext: German Aerospace Center (DLR), Institute of Communications and Navigation, Wessling, Germany 66institutetext: European Southern Observatory (ESO), Karl-Schwarzschild-Straße 2, 85748 Garching bei München, Germany 77institutetext: Department of Physics and Atmospheric Science, Dalhousie University, Halifax, NS, B3H 4R2, Canada 88institutetext: Academia Sinica Institute of Astronomy and Astrophysics (ASIAA), No. 1, Section 4, Roosevelt Road, Taipei 106216, Taiwan 99institutetext: International Centre for Radio Astronomy Research, University of Western Australia, 7 Fairway, Crawley, Perth, Australia 1010institutetext: ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), Australia 1111institutetext: Cosmic Dawn Center (DAWN), Denmark 1212institutetext: Instituto de Astrofísica de Canarias, Vía Láctea, 39020 La Laguna (Tenerife), Spain 1313institutetext: Instituto de Radioastronomía y Astrofísica, Universidad Nacional Autónoma de México, Antigua Carretera a Pátzcuaro # 8701, Ex-Hda. San José de la Huerta, Morelia, Michoacán, México C.P. 58089 1414institutetext: National Radio Astronomy Observatory, Charlottesville, VA, USA 1515institutetext: Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC V6T 1Z1, Canada 1616institutetext: Max–Planck Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany
@@
Massive, intensely star-forming galaxies at high redshift require a supply of molecular gas from their gas reservoirs, replenished by infall from the surrounding circumgalactic medium, to sustain their immense star-formation rates. However, our knowledge of the extent and morphology of their cold-gas reservoirs is still in its infancy. We present the results of stacking 80 hours of JVLA observations of CO(1–0) emission – which traces the cold molecular gas – in nineteen dusty, star-forming galaxies from the AS2VLA survey. The visibility-plane stack reveals extended emission with a half-light radius of kpc, 2–3 more extended than the dust-obscured star formation and more extended than the stellar emission revealed by JWST. Stacking the [C i](1–0) observations for ten galaxies from our parent sample yields a half-light radius 2.6 kpc, marginally smaller than CO(1–0). The CO(1–0) size is also comparable to the [C ii] halos detected around high-redshift star-forming galaxies, suggesting these arise from molecular gas. Photo-dissociation region modelling indicates that the extended CO(1–0) emission arises from clumpy, dense clouds rather than smooth, diffuse gas. Our results show that the bulk (up to 80%) of molecular gas resides outside the star-forming region; with only a small part directly contributing to their current star formation.
@@
The majority of star formation beyond redshift took place in massive, dusty star-forming galaxies (DSFGs111Also know as sub-millimetre galaxies (SMGs).) with star-formation rates (SFRs) of a few 100 – 1000 M⊙ yr-1 (e.g., Magnelli et al. 2011; Dudzevičiūtė et al. 2020; Zavala et al. 2021). While some DSFGs are starbursts triggered by major mergers, many appear to be relatively isolated (Gillman et al., 2024; Hodge et al., 2025). The immense SFRs of these isolated galaxies need to be fed by a continuous supply of cold gas from their extended gas reservoirs and the surrounding circumgalactic medium (CGM, e.g., Tumlinson et al. 2017; Faucher-Giguère & Oh 2023) or by minor mergers with gas-rich satellites (e.g., Narayanan et al. 2015).
@@
In local galaxies, cold, molecular gas is primarily traced by the CO(1–0) rotational line, which has an excitation temperature of just 5.5 K and critical density of cm-3. At high redshift, studies of molecular gas have mostly focused on higher-excitation CO rotational emission lines (see, e.g., Hodge & da Cunha 2020; Tacconi et al. 2020; Walter et al. 2020; Birkin et al. 2021). However, the mid/high-excitation CO emission lines arise from the denser, warmer gas, and are expected to be much more compact than the ground-state CO(1–0) transition (e.g., Leroy et al. 2009; Ivison et al. 2011; Emonts et al. 2016), making them unsuitable for tracing cold, diffuse gas.
@@
In contrast to the higher-excitation CO lines, studies of the ground-state CO(1–0) line in high-redshift galaxies have been scarce and largely focused on the few brightest sources. This is because of the low intrinsic brightness of the CO(1–0) line, which requires a large collecting area and long integrating times to achieve the necessary signal-to-noise ratio (S/N).
@@
The CO(1–0) emission in a number of high-redshift galaxies has been observed by the Green Bank Telescope (e.g., Hainline et al. 2006; Swinbank et al. 2010; Frayer et al. 2011, 2018; Harris et al. 2012; Harrington et al. 2021). However, these spatially unresolved observations do not provide any information about the source size. Early Karl. G. Jansky Very Large Array (JVLA) studies of CO(1–0) emission in bright high-redshift DSFGs (e.g., Ivison et al. 2010, 2011; Riechers et al. 2011b; Thomson et al. 2012) reported half-light radii ranging from a few kpc to 25 kpc (J02399-0136, Thomson et al. 2012; Frayer et al. 2018), often showing multiple CO(1–0) components. However, some of these results have been challenged by deeper observations: for example, Frias Castillo et al. (2022) showed that the supposedly extended CO(1–0) reservoir with high-velocity wings in the galaxy J13120+4242 (Riechers et al., 2011a) is likely an artefact of low signal-to-noise of the original data. To summarise, the sizes of CO(1–0) reservoirs reported by these early studies might be biased towards the brightest (i.e., most extreme) sources, and suffer from low S/N.
@@
More recently, apparently very extended molecular gas reservoirs have been reported in high-density environments of early protoclusters, such as the Spiderweb (, Emonts et al. 2016) which shows a CO(1–0) reservoirs with a half-light radius of kpc CO(1–0), the -kpc CO(4–3) emission around the star-forming galaxy from Ginolfi et al. (2017), or the claimed super-extended, 200 kpc CO(3–2) emission around a quasar cid_346 (Cicone et al., 2021), which was however not confirmed with deeper data (Jones et al., 2023)).
@@
Further evidence for extended gas around high-redshift DSFGs comes from Atacama Large (Sub)Millimetre Array (ALMA) observations of the [C ii] 158-m line. Extended [C ii] “halos” have been observed around high-redshift DSFGs (e.g., Gullberg et al. 2018; Rybak et al. 2019, 2020) as well as around less massive star-forming galaxies (e.g., Fujimoto et al. 2019, 2020; Ginolfi et al. 2020; Ikeda et al. 2025), sometimes on scales as large as 50 kpc.
@@
The ground-state CO(1–0) emission line remains the most robust probe of diffuse molecular gas. In the recent years, there have been systematic surveys of CO(1–0) in high-redshift DSFGs using JVLA (Frias Castillo et al., 2023; Stanley et al., 2023). These have yielded tentative evidence for extended CO(1–0) emission: Frias Castillo et al. (2023) found that the CO(1–0) extracted from moment-0 maps increases with the aperture size out to radii of 5 (40 kpc). Similarly, using 1-resolution JVLA imaging of 14 galaxies from the z-GAL survey (Cox et al., 2023), Stanley et al. (2023) found CO(1–0) reservoirs of up to 25 kpc across. However, as about 50% of the Stanley et al. (2023) sources are gravitationally lensed, the interpretation of the measured source sizes is not straightforward.
In this Paper, we report results of stacking 80 hours of CO(1–0) observations from the recently completed AS2VLA222AS2VLA = ALMA, SCUBA-2, VLA. survey (Frias Castillo et al., 2023). Specifically, using the JVLA in the most compact D-array and aiming at high surface-brightness sensitivity, provides the ideal set-up to study potentially very extended, low surface-brightness cold gas reservoirs around high-redshift DSFGs.
This Paper is structured as follows: in Section 2, we describe our sample selection, details of observations, and the visibility-plane analysis and stacking procedure. We then look for potential cold-gas outflows (Section 3.1), present the ancillary [C i](1–0) data, and use PDR modelling to infer the thermodynamics of extended gas reservoirs (3.3). Finally, we compare our inferred CO(1–0) sizes to observations of dust, [C ii], and stellar emission (Section 4.1 and 4.2) and predictions from hydrodynamical simulations (4.3).
@@ 2.1 Target sample
Our initial sample consists of 30 sub-millimetre selected, star-forming galaxies at from the AS2VLA survey (Frias Castillo et al., 2023, Paper I). The sample is selected from ALMA continuum detections of source taken from 4 deg2 of SCUBA-2 850m maps from the S2CLS (Geach et al., 2017) and S2COSMOS (Simpson et al., 2019) surveys. Individual sources were pin-pointed by high-resolution ALMA and SMA imaging (Stach et al., 2018; Hill et al., 2018; Simpson et al., 2020); the 870-m fluxes range between 7.5 and 18.3 mJy. All sources have spectroscopic redshifts based on mid- CO lines and [C i](1–0) (Birkin et al. 2021; Chen et al. 2022a, b, Chapman et al., in prep.). The sources span stellar mass M⊙ and SFR = 400 – 1700 M⊙ yr-1, with a median SFR of 1010 M⊙ yr-1 (Frias Castillo et al., 2023; Liao et al., 2024). and SFR were derived by Dudzevičiūtė et al. (2020); Simpson et al. (2020) using the high-redshift version of the MagPhys spectral energy distribution modelling code (da Cunha et al., 2015; Battisti et al., 2019). Most of them are on the massive end of the main sequence of star-forming galaxies, but several are offset to the starburst regime (SFR/ than the main-sequence, Frias Castillo et al. 2023). None of the galaxies is strongly gravitationally lensed, as confirmed by ALMA 0.3”-resolution continuum imaging (Stach et al., 2018; Simpson et al., 2020). We can thus directly measure their intrinsic sizes.
@@ 2.2 JVLA observations
The AS2VLA survey used JVLA to observe the CO(1–0) emission (rest-frame frequency: = 115.2712 GHz) in thirty DSFGs at . The surveys comprises of the JVLA programs ID:21A-254 (PI: Hodge) and ID:17A-251 (PI: Walter). The observations were taken between March 2021 and December 2023, under good weather conditions in the D-array configuration. The total time per source including overheads was four hours for almost all sources. The exceptions were AS2COS14.1, which was observed for eight hours, and AS2COS44.1 and AS2COS65.1, which were observed for two hours each.
We observed in the K- and Ka-bands, using a contiguous bandwidth of 2 GHz with 2-MHz spectral channels. Depending on the source redshift, the angular resolution ranged from 2.5 to 5.1 (circularised FWHM); the maximum recoverable scale was 44–60. We manually inspected and calibrated the data using the Common Astronomy Software Applications package (Casa, McMullin et al. 2007) versions 6.4.1 and 6.4.3 and the standard JVLA pipeline.
The observations of the first 17 galaxies (11 robust detections) were presented in Frias Castillo et al. (2023, Paper I). The remaining sources from the completed survey will be presented in a forthcoming paper (Paper III); out of these, an additional eight sources were found to have secure detections of the CO(1-0) emission line. For this work (Paper II), we selected all the sources of the entire sample having secure CO(1-0) detections (peak S/N3 in naturally-weighted cleaned maps of CO(1-0) emission), amounting to a total of 19 sources.
@@ 2.3 Visibility-plane analysis and stacking
We measure the source sizes directly in the visibility-plane. This approach bypasses several issues with the Cleaned images, including varying beam shape, the fact that cleaned images combine the dirty and clean (Gaussian) beam, and the correlated noise in the image-plane. For illustration, we show the stacked moment-0 image in Fig. 1. We achieve an rms sensitivity of 0.014 Jy km s-1 beam-1 with a median beam FWHM 4.2”2.7”, corresponding to K km s-1 pc2 kpc-2 at , the median redshift of our sample. Image-plane deconvolution with the imfit task in Casa does not yield a reliable source size.
@@
For the visibility-plane analysis, we process the data as follows: (1) if the wide-band imaging shows a significant (S/N5) continuum, subtract the continuum (constant signal) from the data using the uvcontsub task; (2) select the velocity range within 1 FWHM from the systemic velocity (using linewidths from mid- CO lines, which are detected at a higher S/N); (3) frequency-average the selected data into a single channel; (4) bin the resulting visibilities in radial bins with a size of 2.5, 5, and 10 k. The fields were centred at the ALMA dust continuum peak (accurate to 0.1). The phase-tracking centre is generally well-aligned with the CO(1–0) emission; we phase-shifted three sources with noticeable offsets (AEG2, CDFN1, CDFN2). For two sources (AS2COS14.1, AS2COS31.1), we found a bright continuum source within the VLA primary beam. We subtracted these companions from the visibilities in Casa, using the ft task.
@@
We fit each binned visibility function with: (1) point-source (constant signal) and (2) an exponential profile. We prefer the exponential profile to the more widely-used Gaussian profile as it provides a more physical description of gas surface brightness distribution in star-forming galaxies 333For the literature source sizes derived using Gaussian profiles, we convert the half-light radius from the Gaussian to exponential profile. (e.g., Leroy et al. 2009; Wang & Lilly 2022) as well as dust and stellar light distribution (e.g., Hodge et al. 2016; Gullberg et al. 2019; Gillman et al. 2024. We utilise Monte-Carlo Markov chain modelling using the emcee package (Foreman-Mackey et al., 2013). We assumed wide, uniform priors in total flux and half-light radius. For each galaxy, we calculate the Akaike and Bayesian information criteria (AIC, BIC) for each model.
Table 1 lists the source sizes inferred from the MCMC modelling. For galaxies for which the extended source is not preferred by both AIC and BIC, we give upper limit on the source sized, based on the 95th percentile from the posterior of the exponential model.
Source ID | CO(1–0) | [C i](1–0) | |||
[arcsec] | [kpc] | [arcsec] | [kpc] | ||
AS2COS1.1 | 4.625 | 2.62 | 17 | 1.36 | 9 |
AS2COS8.1 | 3.581 | 1.05 | 8 | 0.75 | 6 |
AS2COS13.1 | 2.608 | 0.41 | 2.9 | — | — |
AS2COS14.1 | 2.921 | 0.74 | 5.8 | 0.98 | 7.7 |
AS2COS23.1 | 4.341 | 1.89 | 13 | 1.110.44 | 7.63.0 |
AS2COS28.1 | 3.097 | 0.66 | 5.1 | — | — |
AS2COS31.1 | 3.643 | 5.26 | 39 | 0.44 | 3.4 |
AS2COS54.1 | 3.174 | 1.60 | 12 | — | — |
AS2COS65.1 | 2.414 | 0.68 | 5.6 | 0.380.11 | 3.10.9 |
AS2COS66.1a | 3.247 | 2.09 | 17 | — | — |
AS2COS139.1 | 2.219 | 1.08 | 8 | — | — |
AS2UDS10.0 | 3.169 | 3.65 | 28 | — | — |
AS2UDS11.0 | 4.073 | 2.37 | 38 | — | — |
AS2UDS12.0 | 2.520 | 0.80 | 6.6 | 0.400.19 | 3.31.6 |
AS2UDS126.0 | 2.436 | 0.97 | 8.0 | 0.640.28 | 5.32.3 |
AEG2 | 3.668 | 1.43 | 12 | — | — |
CDFN1 | 3.149 | 0.80 | 6 | — | — |
CDFN2 | 4.422 | 1.35 | 9 | — | — |
CDFN8 | 4.144 | 0.89 | 6 | — | — |
AS2COS2.1 | 4.595 | — | — | 2.30 | 15 |
AS2UDS.26 | 3.296 | — | — | 0.67 | 5.1 |
a The [C i] line in AS2COS66.1 falls partially outside the ALMA tuning | |||||
and is not included in this analysis. |